alissak1223:

find the quotient of the complex numbers. express your answer in trigonometric form z1=12(cos(320)+isin(320) z2= 4(cos(80)+isin(80)

snowflake0531:

Which means that we have z1=12(cos(320)+isin(320) z2= 4(cos(80)+isin(80) 12/4 =3 320-80=240 320-80=240 So, if those are the choices, in the link above, it would be D 3(cos(240))+isin(240)

alissak1223:

thank you

surjithayer:

\[z_{1}=12e^{320 \iota}\] \[z _{2}=4 e^{80 \iota}\] \[\frac{ z_{1} }{ z_{2 }}=\frac{ 12 e ^{320 \iota} }{ 4 e^{80 \iota } }=3 e^{320 \ \iota-80 \ \iota}=3e^{240 \ \iota }\] \[=3 (cos 240 +\iota sin 240)\]